# Energy dynamics in a class of local random matrix Hamiltonians

### v2

@@ II.4 Out of time ordered correlator
$\braket{0}{V(t)V(0)V(t)V(0)}{0}\\ =\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{kq}\braket{k}{V(t)V(0)}{q}\braket{q}{V(t)V(0)}{k}.$ (43)
@@
$\braket{k}{V(t)V(0)}{q}=\delta_{kq}(\cos^{2}(k)+\sin^{2}(k)e^{-2i\epsilon_{k}t})\\ +i\delta_{k,q+\pi}\cos(k)\sin(k)(1-e^{-2i\epsilon_{k}t})$ (44)
@@ II.5 Finite temperature
$C_{\beta}(t)=Z_{1}^{-1}\operatorname{tr}_{1}[e^{-\beta\Delta},V(t)V(0)]$ (47)
@@ III Arbitrary number of mutually non-commuting terms
In the previous section, we proved that irreducible words of two terms vanish in the large $q$ limit, when arranged so that they overlap on a single site. At the same time, for the model in Fig. 4(b), the irreducible words of those terms provably vanish by a standard result in free probability [we will make those connections more explicit in Sec. V]. Based on these facts, as well as empirical evidence that this also holds for the terms arranged as in sketch (a), we conjecture that in all scenarios presented in Fig. 4, the irreducible words of the terms vanish. This gives way to a similar single-particle mapping as developed in Sec. II, but here the particle hops on the Bethe lattice with coordination number $z$. We study in a bit more detail the particular case of $z=3$ with the terms arranged as in panel (a) of that figure, confirming that numerically, the single particle hopping on the Bethe lattice gives the empirically correct density of states and energy-energy correlator.
Figure 4: Various few-body scenarios of HI local terms interacting. Under a conjecture that irreducible words always vanish, (a),(b),(c) all map to a single particle hopping on the Bethe lattice with coordination number $z=3$ (for scenario (b) the irreducible words provably vanish). Sketch (a) can be viewed as a short 1D chain in PBC. Sketch (b) is strictly zero dimensional with all $z=3$ terms acting on the same pair of sites. Sketches (c) and (d) can be viewed as parts of a larger lattice of local terms interacting at a site in 2D, and (d) maps to the $z=4$ Bethe lattice again under the conjecture that irreducible words of the terms vanish in the large $q$ limit.
@@ III.1 Hopping on the Bethe lattice
etc. Reference [50] points out, however, that in the |$\ket{d}$ basis, the hopping matrix $\Delta$ becomes simple:
$\Delta\ket{0}=\sqrt{z}\ket{1}$
$\Delta\ket{1}=\sqrt{z}\ket{0}+\sqrt{z-1}\ket{2}$
$\Delta\ket{d}=\sqrt{z-1}\big{[}\ket{d-1}+\ket{d+1}\big{]}\quad d\geq 2$
@@ III.2 Density of states
where the set $[z]=\{0,1,2,\dots,z-1\}$ and $h_{b}$ is again shorthand for a correlator defined by the word $b$. Under the assumption that the irreducible correlators vanish in a large $q$ limit, we obtain the correspondence
$\braket{H^{n}}=\braket{0}{\Delta^{n}}{0},$ (58)
@@
Indeed Fig. 5(c) compares this equation with $z=3$ to the ED numerics for three terms in PBC with $q=12$. Their similarity is consistent with the assumption that all irreducible correlators vanish in the large $q$ limit.
@@ III.3 Energy dynamics
where $h$ stands for any one of the $h_{i}$ in the Hamiltonian. In a double series expansion, i.e. Eq. (28), the combinatorial problem is again to calculate $\braket{H^{n}hH^{m}h}$ in the large $q$ limit. Under our conjecture, only the reducible correlators survive, and we thus need to count the number of words of the form $aibi$ with $a\in[z]^{n}$ and $b\in[z]^{m}$, for some fixed $i$, that are fully reducible. To count these, we require that $a$ reduces to some irreducible word, say $c$, (equivalently a site $c$ on the Bethe lattice) and $ibi$ reduces to the inverse of that word, $c^{-1}$. Then, we sum over all points $c$. We show the details of this counting problem in Appendix A and give the result here.
@@
$W_{n}^{d}=\frac{\braket{d}{\Delta^{n}}{0}}{\sqrt{P_{d}}}$ (62)
since $\braket{d}{\Delta^{n}}{0}$ counts the number of walks that end at distance $d$ up to a normalization. We can identify a non-local operator $X$ such that  [^3] We were not able to identify an operator $V$ such that $C(t)=\braket{0}{V(t)V(0)}{0}$ as we were in the $z=2$ case, the main issue being that on the Bethe lattice ($z>2$), points are not labeled by $d$ anymore.
$C(t)=\braket{0}{X(t)}{0}.$ (63)
This operator can be read off from Eq. (61) and is best understood in matrix form. In the |$\ket{d}$ basis, we have
@@
and $C(t)$ can then be understood as the following experiment: a particle begins at the origin at time $-t$. Then, at time 00, the system is kicked with the operator $X$ everywhere on the lattice; then $C(t)$ is the amplitude for the particle to return to the origin at time $t$.
Since momentum is no longer a good quantum number, an analytical calculation become more difficult. We can, however, run the hopping dynamics numerically on a large but finite Bethe lattice of radius $R$. Introducing a finite-size cutoff means the dynamics will deviate from the infinite lattice case after a certain time $t^{*}=(2\sqrt{z-1})^{-1}R$, which is the timescale for the particle to propagate ballistically to the edge. In Fig. 5(c) we compute $C(t)=\braket{0}{X(t)}{0}$ for the $z=3$ Bethe lattice for $t<t^{*}$ and compare it to the numerical solution for three random matrix terms arranged as in Fig. 5(a). Here, again a single realization is shown, and we found it necessary to show a very small $q=6$ to observe noticeable error. For $q=10$, the curves were essentially indistinguishable. In the inset, we also show one standard deviation from the average as a function of time, which is small (roughly $1.5\%$ of the average) and flat on the plotted timescale.
@@
In this totally non-commuting scenario, by writing $H=h+\tilde{h}$, where $\tilde{h}$ stands for all other $z-1$ terms in $H$, we can conjecture that, physically, $z\geq 3$ decays faster than $z=2$ because in the former case, the “bath” Hamiltonian $\tilde{h}$ is chaotic, while for $z=2$ it is not.
@@ IV Energy transport in the extended chain
We emphasize that in the diffusion equation, $r$ refers to a site whereas in the chain $r$ refers to a bond. So in an OBC chain of length $L$, which has $L-1$ bonds, the rightmost site is labeled $r=L-2$. Since the diffusion equation is linear in the density, it is easily solved numerically by treating the equation as an imaginary time Schrodinger equation and applying the propagator to an initial configuration; in our case the initial condition being $\rho_{0}(0)=1$ and $\rho_{r}(0)=0$ for $r\geq 1$.
@@
$\braket{\psi}{A}{\psi}=\operatorname{tr}[A]+O\bigg{(}\frac{\operatorname{tr}[A^{\dagger}A]^{1/2}}{q^{L/2}}\bigg{)}$ (71)
as $L\rightarrow\infty$ for a single realization of a Haar random state |$\ket{\psi}$ of $L$ spins. One can derive this formula using the Weingarten calculus, but the idea of using a random vector to estimate traces goes all the way back to [54]. With this method, one simply needs to calculate the two vectors $h_{r}e^{-iHt}\ket{\psi}$ and $e^{-iHt}h_{0}\ket{\psi}$, so that $C_{r}(t)$ is given by their overlap. While one or only a handful of realizations usually suffice in this context, since we must average over many RMT realizations anyways, we also average over a new Haar random state for each run, which can only improve the accuracy. We evolve the states using Krylov time evolution with a basis dimension of 12 [55].
@@ V Spectra and Free probability
Figure 7: Sketches of two small “local” HI random matrix Hamiltonians where a free convolution can be readily applied to compute the density of states. Those are calculated numerically and shown, for (a) and (b), in Fig. 8 (a) and (b), respectively.
@@
That the two terms are asymptotically free is consistent with our earlier results, as the free convolution of two symmetric Bernoulli distributions is the arcsine distribution [23]. In fact, considering $z$ mutually non-commuting HI terms, the conjecture we stated in Sec. III that irreducible words vanish, is also equivalent to the statement that those terms are asymptotically free by an extension of the above argument for two terms.
Once true locality of interaction is introduced into an RMT, however, these tools do not directly apply. While the Hamiltonian will be of the form $A+B+\cdots$, the individual terms are generally no longer free, since their eigenvectors are no longer in generic position. Rather, since most of the terms commute, they actually share eigenvectors, and, if they are drawn independently, then they behave as classically independent random variables. A few works have begun to address this question of the density of states of a local random matrix model in the large $q$ limit, and in particular how to address the mixture of free and classical independence [21, 20, 22]. In the mathematical literature the topic goes by the name $\Lambda-$freeness or $\varepsilon-$freeness [56, 57]. In particular, in [58], their “XY-model” is identical in structure to our local nearest-neighbor chain, but they draw the local terms from the GUE and a bound on the maximum eigenvalue is derived, but these works are still inconclusive insofar as producing a concrete example of the density of states of a local random matrix theory.
On the other hand, based on intuition from a non-interacting system of $L$ spins or numerical exact diagonalization of interacting spin chains, one expects that if first the $L\rightarrow\infty$ limit is taken, then the density of states should be a Gaussian. In fact, this was proven for local translation-invariant random $q=2$ spin chains [59] and for disordered $q=2$ “spin-glasses” on graphs with bounded degree in [19]. In Appendix C we also give a heuristic argument for a Gaussian density of states in a 1D chain which is independent of the microscopic details, in particular independent of $q$. For the local HI random matrix, one can see in Fig. 8(d) that a modest $L=14$ chain with small $q=2$ has an approximately Gaussian density of states of variance $\sigma^{2}=L$. It is still an interesting question as to how precisely how this non-interacting character of the density of states emerges as $L$ grows, and in particular if the infinite $q$ limit is taken first, as in the free probability literature.
@@
This is consistent with the mathematical literature on mixtures of free and classical independence [22] where it is first noted that the situation cannot be split into two groups. These chain configurations with more than two terms might serve as non-trivial examples of asymptotic $\varepsilon$-independence. In particular, our conjecture about the mixed correlators of the three sites vanishing is equivalent to asymptotic $\epsilon$-freeness. As mentioned before, the development of this framework is in its early stages and there do not seem to be any results explicitly computing the density of states. It may be possible to develop a numerical $\epsilon$-free convolution similar to the subordination iteration method.
@@ VI Conclusion and Outlook
In this paper we have studied the evolution of the local energy density under the dynamics of a local random matrix Hamiltonian whose bond operators square to one. This simplification allowed for the analytical calculation of the density of states, a two-point function, and an OTOC for the two-term and large $q$ scenario (Sec. II). Under a conjecture that irreducible correlators still vanish, we found a similar mapping of a generalized model with $z$ non-commuting terms (Sec. III) onto walks on the Bethe lattice and we calculated energy-energy correlators, finding that the two point function always decays as $t^{-3}$ for any $z\geq 3$. We also examined energy transport in the small $q$ limit and with increasing $L$, finding evidence of a slow convergence towards diffusion, but the data are insufficient to conclude that energy transport in this model falls within the universality class of diffusion. We then moved on to study the simpler question of the density of states in Sec. V. We discussed connections to free probability theory, and showed numerically that in some few-body cases, a classical convolution of the Hamiltonian terms followed by a free convolution can yield the correct density of states despite the technical issue of some terms commuting and some not, but that this method fails in general. We conclude with a few comments and ideas moving forward.
It is interesting that the density of states for the $L=4$ PBC chain [Fig. 8(a)] has a peak in the middle of the band which is reminiscent of the van-Hove singularity for single-particle hopping in 2D [47], while the $L=6$ PBC chain [Fig. 8(c)] also has the qualitative shape of the density of states for hopping in 3D, i.e. a relatively flat central region. Supposing that that irreducible words of the local terms still vanish after the introduction of commutativity, we can speculate that the density of states for $L$ terms in PBC could be derived by considering hopping on some kind of $L/2$ dimensional hypercube-like lattice. This is consistent with expectations in the sense that as we take the dimensionality of the hopping to infinity, the density of states converges to a Gaussian (via the central limit theorem).
@@
and compared against the standard expression for $P(r)$ derived in [62]. To further diagnose chaos, we need to calculate at least two-trace quantities such as the spectral form factor (SFF). Is there a mapping to the hopping picture for two-trace quantities like the SFF, at least for the two-term model? Such a calculation should involve counting words that are not fully reducible, and will be the subject of future work.

### v3

@@ II.4 Out of time ordered correlator
$\braket{0|V(t)V(0)V(t)V(0)|0}\\ =\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{kq}\braket{k|V(t)V(0)|q}\braket{q|V(t)V(0)|k}.$ (43)
@@
$\braket{k|V(t)V(0)|q}=\delta_{kq}(\cos^{2}(k)+\sin^{2}(k)e^{-2i\epsilon_{k}t})\\ +i\delta_{k,q+\pi}\cos(k)\sin(k)(1-e^{-2i\epsilon_{k}t})$ (44)
@@ II.5 Finite temperature
$C_{\beta}(t)=Z_{1}^{-1}\operatorname{tr}_{1}[e^{-\beta\Delta}V(t)V(0)]$ (47)
@@ III Arbitrary number of mutually non-commuting terms
In the previous section, we proved that irreducible words of two terms vanish in the large $q$ limit, when arranged so that they overlap on a single site. This result continues to hold when considering $z$ terms which are mutually non-commuting and arranged with all bonds overlapping on at least a single site, which follows from Refs. [26, 51]. This observation gives way to a similar single-particle mapping as developed in Sec. II, but here the particle hops on the Bethe lattice with coordination number $z$. We study in a bit more detail the particular case of $z=3$ with the terms arranged as in Fig. 4 panel (a), confirming that numerically, the single particle hopping on the Bethe lattice gives the empirically correct density of states and energy-energy correlator.
Figure 4: Various few-body scenarios of HI local terms interacting. (a),(b),(c) all map to a single particle hopping on the Bethe lattice with coordination number $z=3$ (for scenario (b) the irreducible words provably vanish). Sketch (a) can be viewed as a short 1D chain in PBC. Sketch (b) is strictly zero dimensional with all $z=3$ terms acting on the same pair of sites. Sketches (c) and (d) can be viewed as parts of a larger lattice of local terms interacting at a site in 2D, and (d) maps to the $z=4$ Bethe lattice.
@@ III.1 Hopping on the Bethe lattice
The central idea of this section is that the combinatorics of reducible correlators of $z$ HI terms finds a natural realization as walks on a Bethe lattice of coordination number $z$ with the $z=2$ case being the 1D walks studied in the previous section. With respect to some fixed origin, we label positions on the Bethe lattice by irreducible words $c$ of $z$ symbols; say $0,1,\dots,z-1$. We denote a single-particle configuration with a particle living on site $c$ by $|c)$. The $z=3$ Bethe lattice is sketched in Fig. 5(b). The origin is the empty word $\emptyset$, and the points at “radius” $1$ (the first dashed ring) are the words $0,1,2$ and so on.
@@
[…]
Let $\Delta$ be the NN hopping Hamiltonian on the Bethe lattice. The Hamiltonian acts on the $|c)$ states by uniformly hopping to all nearest neighbor points, e.g. for $z=3$ its action is
@@
[…]
@@ III.2 Density of states
where the set $z]=\{0,1,2,\dots,z-1\}$ and $h_{b}$ is again shorthand for a correlator defined by the word $b$. Under the assumption that the irreducible correlators vanish in a large $q$ limit, we obtain the correspondence
$\braket{H^{n}}=\braket{0|\Delta^{n}|0},$ (58)
@@
Indeed Fig. 5(c) compares this equation with $z=3$ to the ED numerics for three terms in PBC with $q=12$, the result being consistent with irreducible correlators vanishing in the large $q$ limit.
@@ III.3 Energy dynamics
where $h$ stands for any one of the $h_{i}$ in the Hamiltonian. In a double series expansion, i.e. Eq. (28), the combinatorial problem is again to calculate $\braket{H^{n}hH^{m}h}$ in the large $q$ limit. Only the reducible correlators survive, and we thus need to count the number of words of the form aibiaibi with $a\in[z]^{n}$ and $b\in[z]^{m}$, for some fixed $i$, that are fully reducible. To count these, we require that $a$ reduces to some irreducible word, say $c$, (equivalently a site $c$ on the Bethe lattice) and ibiibi reduces to the inverse of that word, $c^{-1}$. Then, we sum over all points $c$. We show the details of this counting problem in Appendix A and give the result here.
@@
$W_{n}^{d}=\frac{\braket{d|\Delta^{n}|0}}{\sqrt{P_{d}}}$ (62)
since $\braket{d|\Delta^{n}|0}$ counts the number of walks that end at distance $d$ up to a normalization. We can identify a non-local operator $X$ such that  [^3] We were not able to identify an operator $V$ such that $C(t)=\braket{0|V(t)V(0)|0}$ as we were in the $z=2$ case, the main issue being that on the Bethe lattice ($z>2$), points are not labeled by $d$ anymore.
$C(t)=\braket{0|X(t)|0}.$ (63)
[…]
@@
and $C(t)$ can then be understood as the following experiment: a particle begins at the origin at time t-t. Then, at time 0, the system is kicked with the operator $X$ everywhere on the lattice; then $C(t)$ is the amplitude for the particle to return to the origin at time $t$.
Since momentum is no longer a good quantum number, an analytical calculation becomes more difficult. We can, however, run the hopping dynamics numerically on a large but finite Bethe lattice of radius $R$. Introducing a finite-size cutoff means the dynamics will deviate from the infinite lattice case after a certain time $t^{*}=(2\sqrt{z-1})^{-1}R$, which is the timescale for the particle to propagate ballistically to the edge. In Fig. 5(c) we compute $C(t)=\braket{0|X(t)|0}$ for the $z=3$ Bethe lattice for $t<t^{*}$ and compare it to the numerical solution for three random matrix terms arranged as in Fig. 5(a). Here, again a single realization is shown, and we found it necessary to show a very small $q=6$ to observe noticeable error. For $q=10$, the curves were essentially indistinguishable. In the inset, we also show one standard deviation from the average as a function of time, which is small (roughly $1.5\%$ of the average) and flat on the plotted timescale.
@@
In this totally non-commuting scenario, by writing $H=h+\tilde{h}$, where $\tilde{h}$ stands for all other z1z-1 terms in $H$, we can conjecture that, physically, $z\geq 3$ decays faster than $z=2$ because in the former case, the “bath” Hamiltonian $\tilde{h}$ is chaotic, while for $z=2$ it is not.
@@ IV Energy transport in the extended chain
We emphasize that in the diffusion equation, $r$ refers to a site whereas in the chain $r$ refers to a bond. So in an OBC chain of length $L$, which has L1L-1 bonds, the rightmost site is labeled r=L2r=L-2. Since the diffusion equation is linear in the density, it is easily solved numerically by treating the equation as an imaginary time Schrodinger equation and applying the propagator to an initial configuration; in our case the initial condition being $\rho_{0}(0)=1$ and $\rho_{r}(0)=0$ for $r\geq 1$.
@@
$\braket{\psi|A|\psi}=\operatorname{tr}[A]+O\bigg{(}\frac{\operatorname{tr}[A^{\dagger}A]^{1/2}}{q^{L/2}}\bigg{)}$ (71)
[…]
@@ V Spectra and Free probability
Figure 7: Sketches of two small local HI random matrix Hamiltonians where a free convolution can be readily applied to compute the density of states. Those are calculated numerically and shown, for (a) and (b), in Fig. 8 (a) and (b), respectively.
@@
That the two terms are asymptotically free is consistent with our earlier results, as the free convolution of two symmetric Bernoulli distributions is the arcsine distribution [28]. In fact, considering $z$ mutually non-commuting HI terms (Sec. III), the statement that irreducible words vanish is also equivalent to the statement that those terms are asymptotically free by an extension of the above argument for two terms.
Once true locality of interaction is introduced into an RMT, however, these tools do not directly apply. While the Hamiltonian will be of the form $A+B+\cdots$, the individual terms are generally no longer free, since their eigenvectors are no longer in generic position. Rather, since most of the terms commute, they actually share eigenvectors, and, if they are drawn independently, then they behave as classically independent random variables. A few works have begun to address this question of the density of states of a local random matrix model in the large $q$ limit, and in particular how to address the mixture of free and classical independence [26, 25, 27]. In the mathematical literature the topic goes by the name $\Lambda$-freeness or $\varepsilon$-freeness [64, 65]. In particular, in [26], their “XY-model” is identical in structure to our local nearest-neighbor chain, but these works are still inconclusive insofar as producing a concrete example of the density of states of a local random matrix theory.
On the other hand, based on intuition from a non-interacting system of $L$ spins or numerical exact diagonalization of interacting spin chains, one expects that if first the $L\rightarrow\infty$ limit is taken, then the density of states should approach a Gaussian distribution. In fact, a Gaussian density of states was proven for local translation-invariant random $q=2$ spin chains [66] and for disordered $q=2$ “spin-glasses” on graphs with bounded degree in [24]. In Appendix C we also give a heuristic argument for the 1D chain having Gaussian moments, which is independent of the microscopic details and in particular independent of $q$. For the local HI random matrix, one can see in Fig. 8(d) that a modest $L=14$ chain with small $q=2$ has an approximately Gaussian density of states of variance $\sigma^{2}=L$. It is still an interesting question as to precisely how this non-interacting character of the density of states emerges as $L$ grows for a particular model, and in particular if the infinite $q$ limit is taken first, as in the free probability literature.
@@
This is consistent with the mathematical literature on mixtures of free and classical independence [27] where it is first noted that the situation cannot be split into two groups. These chain configurations with more than two terms serve as non-trivial examples of asymptotic $\varepsilon$-independence. As mentioned before, however, the development of this framework is in its early stages and there do not seem to be any results explicitly computing the density of states. It may be possible to develop a numerical $\epsilon$-free convolution similar to the subordination iteration method.
@@ VI Conclusion and Outlook
In this paper we have studied the evolution of the local energy density under the dynamics of a local random matrix Hamiltonian whose bond operators square to one. This simplification allowed for the analytical calculation of the density of states, a two-point function, and an OTOC for the two-term and large $q$ scenario (Sec. II). We found a similar mapping of a generalized model with $z$ non-commuting terms (Sec. III) onto walks on the Bethe lattice and we calculated energy-energy correlators, finding that the two point function always decays as $t^{-3}$ for any $z\geq 3$. We also examined energy transport in the small $q$ limit and with increasing $L$, finding evidence of a slow convergence towards diffusion, but the data are insufficient to conclude that energy transport in this model falls within the universality class of diffusion. We then moved on to study the simpler question of the density of states in Sec. V. We discussed connections to free probability theory, and showed numerically that in some few-body cases, a classical convolution of the Hamiltonian terms followed by a free convolution can yield the correct density of states despite the technical issue of some terms commuting and some not, but that this method fails in general. We conclude with a few comments and ideas moving forward.
Looking towards the full 1D local HI RMT model, the fact that only reducible words survive the large $q$ limit will be crucial to its solution. That only the fully reducible words survive in general, even when some terms commute and some do not, follows from recent work in $\varepsilon$-free probability theory [26, 51] as applied to our simplified situation where $h_{i}^{2}=1$. To confirm this on a numerical level, in Appendix. D we look explicitly at $L=4$ and $L=6$ PBC HI RMT chains and find that only the reducible words contribute to the moments as $q\rightarrow\infty$.
With this in mind, a mapping to a particle hopping on more general graphs may be obtained. It is interesting that the density of states for the $L=4$ PBC chain [Fig. 8(a)] has a peak in the middle of the band which is reminiscent of the van-Hove singularity for single-particle hopping in 2D [55], while the $L=6$ PBC chain [Fig. 8(c)] also has the qualitative shape of the density of states for hopping in 3D, i.e. a relatively flat central region. We can speculate that the density of states for $L$ terms in PBC could be derived by considering hopping on some kind of $L/2$ dimensional hypercube-like lattice, where the trajectories on this lattice compute the combinatorics of reducible words. This picture is consistent with expectations in the sense that as we take the dimensionality of the hopping to infinity, the density of states approaches a Gaussian.
@@
and compared against the standard expression for $P(r)$ derived in [69]. To further diagnose chaos, we need to calculate multi-trace quantities such as two-point and higher-point spectral form factors (SFFs). Is there a mapping to the hopping picture for two-trace quantities like the two-point SFF, at least for the two-term model? Such a calculation should involve counting words that are not fully reducible, and will be the subject of future work.